OpenStudy (anonymous):

x equals quantity of negative 7 plus or minus I square root of 6 all over 2 x equals quantity of negative 7 plus or minus I square root of 3 all over 2 x equals quantity of 7 plus or minus I square root of 6 all over 2 x equals quantity of 7 plus or minus I square root of 3 all over 2

OpenStudy (anonymous):

Question: Solve x2 - 7x = -13.

OpenStudy (anonymous):

The first post is the answers, sorry I had messed up

OpenStudy (anonymous):

start with \[x^2-7x+13=0\] which looks like \[ax^2+bx+c=0\]with \[a=1,b=-7,c=13\]

OpenStudy (anonymous):

then the solution is \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with the \(a,b,c\) as above

OpenStudy (anonymous):

I don't understand how I would get the answer from that? @satellite73

OpenStudy (anonymous):

this is the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

where you see an \(a\) put \(1\) where you see \(b\) put \(-7\) where you see \(c\) put \(13\)

OpenStudy (anonymous):

that is the only way to get the answer

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

B is a letter

OpenStudy (anonymous):

In my multiple choices I mean lol

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-(-7)\pm\sqrt{(-7)^2-4\times 1\times 13}}{2\times 1}\]

OpenStudy (anonymous):

so no definitely not B since \(-(-7)=7\)

OpenStudy (anonymous):

Im thinking C then?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73